# Auto-Calibration and Biconvex Compressive Sensing with Applications to Parallel MRI

### v2

@@ 4.2 A sufficient condition
The derivation of the lemma is a direct application of Lemma 7 of [flinth_2017] which relies on the sub-differential of $\partial(\|\cdot\|_{(1,2)})$. Given $\textbf{Z}\in\mathbb{C}^{kN\times C}$, the sub-differential $\partial_{\textbf{Z}}(\|\cdot\|_{(1,2)})$ is the set
$\{\textbf{V}\in\mathbb{C}^{kN\times C}|\textbf{V}_{T_{i}}:=\frac{\textbf{Z}_{T_{i}\times[C]}}{\|\textbf{Z}_{T_{i}\times[C]}\|_{F}},\text{if }\|\textbf{Z}_{T_{i}\times[C]}\|_{F}\neq 0,\|\textbf{V}_{T_{i}}\|_{F}\leq 1,otherwise\}.$
In the following sections, we will prove that with the assumptions we have made, the parameters $\delta$ and $\beta$ will be small with high probability. We will also construct a dual certificate $\boldsymbol{\mathcal{Y}}$ with small $\eta$, $\theta$ and $\tau$. Most of the proof procedure follows well-established methods in the compressive sensing literature, such as those described in [ling_strohmer_2015] and [flinth_2017].
@@ 4.3 Conditioning of $\mathbf{A}_{S}$
Next, we give the conditions on when our measurement matrix $\mathbf{A}$ satisfies the condition $\|\textbf{A}_{S}^{*}\textbf{A}_{S}-\textbf{I}_{S}\|\leq\delta$. Recall our forward matrix $\mathbf{A}_{S}=\mathbf{\mathbf{F}}_{\Omega}\mathbf{\Phi}_{S}$, the challenges come from the fact that its rows are dependent. Instead of bounding the norm directly, we apply the triangular inequality and work on bounding the two terms in $\underbrace{\|\mathbf{\Phi}_{S}^{*}\mathbf{\textbf{F}}_{\Omega}^{*}\mathbf{\textbf{F}}_{\Omega}\mathbf{\Phi}_{S}-\mathbf{\Phi}_{S}^{*}\mathbf{\Phi}_{S}\|}_{(1)}+\underbrace{\|\mathbf{\Phi}_{S}^{*}\mathbf{\Phi}_{S}-\mathbf{I}_{S}\|}_{(2)}$ separately. Indeed, even when the full Fourier space is sampled, i.e., $\mathbf{F}_{\Omega}=\mathbf{F}$, the first term (1) is zero and we would at the very least, require the condition (2) to be satisfied.
@@
$\|\mathbf{\Phi}_{S}^{*}\mathbf{\Phi}_{S}-\textbf{I}_{S}\|\leq\eta$ (35)
@@
The probability is over the randomness in B. The proof of the Lemma is derived from Lemma 4.3 of [ling_strohmer_2015], where the authors consider the case B to be fixed but the rows of $\mathbf{\Psi}_{S}$ are chosen uniformly at random with replacement from the DFT matrix.
@@
The next lemma is concerned with the tail bound for $\|\mathbf{\Phi}_{S}^{*}\mathbf{\textbf{F}}_{\Omega}^{*}\mathbf{\textbf{F}}_{\Omega}\mathbf{\Phi}_{S}-\mathbf{\Phi}_{S}^{*}\mathbf{\Phi}_{S}\|$.
@@
Suppose that the rows of $\mathbf{F}_{\Omega}$ are chosen uniformly at random without replacement from $\mathbf{F}$, where $\mathbf{F^{*}F}=\mathbf{F}\mathbf{F}^{*}=\frac{N}{L}$. For a realization of $\mathbf{\Phi}_{S}$ with $\|\mathbf{\Phi}_{S}^{*}\mathbf{\Phi}_{S}-\mathbf{I}_{S}\|\leq\eta_{1}$,
[…]
, provided,
@@ 4.4 Dual Certificate
Suppose that either each column of H is chosen independently and uniformly at random with replacement from the columns of an orthonormal basis $\textbf{W}\in\mathbb{C}^{k\times k}$ or that each column of H is chosen independently and uniformly at random from the complex sphere $S_{\mathbb{C}}^{k-1}$ (e.g., $h_{l}\in\mathbb{C}^{k}\text{ with }\|h_{l}\|=1$). Conditioned on $\|\textbf{A}_{S}^{*}\textbf{A}_{S}-\textbf{I}\|\leq\eta$, for any $s>0$, if
@@
Remark: Note that increasing the number of channels $C$ increases the probability of recovery. Furthermore, the condition on $L$ is automatically met due to the bound $L\sim O(kn\log(k^{2}n^{2}N))$ from Lemma 5.
Equivalently, we can state the result in Lemma 8 as $\max\limits_{j\in[|S^{c}|]}\|\mathcal{P}_{S^{c}_{j}}\boldsymbol{\mathcal{Y}}\|_{F}<1/2$ with probability at least $1-N^{-\alpha}$ if
@@
where $\tilde{C}_{\alpha}$ grows linearly with respect to $\alpha$.
@@ 5.1 Empirical success rate for the proposed method
For the first set of experiments, we follow the same setup as in Ling and Strohmer [ling_strohmer_2015] and calculate the empirical probability of success of the proposed method. Fix the number of measurements $L=128$ and the underlying signal dimension $N=256$, and solve the optimization problem (28) for varying $(k,n),k,n\in[15]$ and $C\in\{1,2,4\}$. For each $C$ and $(k,n)$, we generate $\textbf{z}\in\mathbb{C}^{N}$ with random support with cardinality $n$ and the nonzero entries of z and $\textbf{H}\in\mathbb{C}^{k\times C}$ drawn from $\mathcal{N}(0,1)$. Each time, we draw $L$ points uniformly from the Fourier space. We repeat the experiment ten times, and each time new H and z are generated.
We count an experiment as a success if the relative error between the solution of our optimization problem $\hat{\textbf{X}}$ and the truth $\textbf{X}_{0}$ is less than $0.01\%$, i.e., $\frac{\|\hat{\textbf{X}}-\textbf{X}_{0}\|_{F}}{\|\textbf{X}_{0}\|_{F}}\leq 10^{-4}$. And the empirical success rate is a fraction defined as (total number of successes) / (total number of experiments). We use the cvxpy [agrawal2018rewriting, diamond2016cvxpy] package of Python with the MOSEK [mosek] solver to solve the mixed norm optimization problem (28).
@@
Next, we illustrate the barrier for reconstruction with the proposed framework when $\boldsymbol{\Psi}$ is a Wavelet transform. From the theory, a small mutual coherence, i.e., $\mu^{\Psi}_{max}\sim O(1)$, is required for the performance of the compressive sensing method. However, the Daubechies wavelet, which can be used to sparsify MRI images [Chun_Adcock_Talavage_2014], has high coherence. We followed the non-uniform sampling scheme in the Fourier measurements as suggested by [Chun_Adcock_Talavage_2014, ADCOCK_HANSEN_POON_ROMAN_2017] as an attempt to improve the results. However, even when $L=N$ (when Fourier space is fully sampled), recovery is poor. We use the Daubechies 4 wavelet as an illustration, and the results are shown in Figure 4.
@@ 5.1.1 Improved results compared to $\ell_{1}$ minimization
5.1.2 Minimal L required for exact recovery is proportional to $kn$
@@ 5.2 Reconstruction of the analytical phantom
We use the analytical phantom [Guerquin-Kern_Lejeune_Pruessmann_Unser_2012] of size $256$ by $256$ and randomly generated coil sensitivities following the Biot–Savart law. It is important to note that we do not manually enforce sparsity in our signal. For example, one may manually use a sparse estimation of the phantom x, e.g., by crafting the signal by thresholding $\textbf{z}=\boldsymbol{\Psi}\textbf{x}=0,if|\boldsymbol{\Psi}\textbf{x}|<\tau$ and set $\textbf{x}=\boldsymbol{\Psi}^{*}\textbf{z}$. The thresholding effectively reduces the dimensionality of the problem and would favor the reconstruction results. However, such manipulation is only possible in simulations when the true signal is known beforehand. Our approach is more flexible and allows for sparsity deficiency in the underlying signal.
@@
To assess the performance of the proposed method with respect to the number of coils, we plot the average reconstruction error versus the number of measurements per coil, for different numbers of coils and levels of Gaussian noise. We adopted a similar setup as that in [otazo2009distributed, Otazo_Kim_Axel_Sodickson_2010] where the multi-coil signal was generated by multiplying a small analytical phantom (N = 529) by the coil sensitivities (k=6) and optionally adding Gaussian noise. The average reconstruction error (relative $\ell_{2}$ distance) over 10 different random Fourier space undersampling realizations was computed for different numbers of measurements and coil elements. A DCT transform was employed to sparsify the image and no strict sparsity was enforced in the signal. The number of required measurements to achieve the noise floor reconstruction decreases with the number of coils for both the noise-free and the noisy case. These trends are consistent with previous numerical experiments for the known calibration pMRI reconstruction, e.g. [otazo2009distributed]. The plots confirm that with a large number of coils, the number of measurements per coil sufficient to achieve a baseline level of reconstruction error is significantly smaller than the true dimensionality of the signal.
@@ 5.3 Comparison with the ENLIVE method
Next, we compare our reconstruction results with the ENLIVE formulation (17) in the paper by Holme et al. [Holme_Rosenzweig_Ong_Wilke_Lustig_Uecker_2019]. We reproduced the reconstructed brain and knee images exactly as shown in the original paper using the data and source code provided in https://github.com/mrirecon/enlive. We then compared these results with our own reconstructions. Figures 11 and 12 show calibrationless variable-density Poisson-disc undersampled reconstructions with differing undersampling factors comparing ENLIVE to our proposed method. Figure 13 shows the reconstructions of the simulated analytical phantom from uniformly subsampled Fourier space with different undersampling factors. As a reference, the corresponding sampling patterns are shown in the third columns of each of the plots.
@@ 6.1 Proof of Lemma 3
Recall that $\boldsymbol{\Phi}=\begin{bmatrix}(\bf{\bar{q}}_{1}\bigotimes\bf{b}_{1})^{*}\\ \vdots\\ \bf{\bar{q}}_{N}\bigotimes\mathbf{b}_{N})^{*}\end{bmatrix}$, where each row of $\boldsymbol{\Phi}$ depends only on one row of B and $\boldsymbol{\Psi}^{*}$. Due to symmetry, the probability distribution of $\mathbb{\Phi}$ is unchanged if we assume B is fixed but the rows of $\mathbf{\Psi}^{*}$ are random. More precisely, we assume the rows of $\mathbf{\Psi}^{*}\in\mathbb{C}^{N\times M}$ are chosen uniformly at random without replacement from some $\mathbf{\Psi}_{0}^{*}\in\mathbb{C}^{N\times M}$ with $\mathbf{\Psi}_{0}^{*}\mathbf{\Psi}_{0}=\textbf{I}_{N}$. Under such a model, the rows of $\mathbf{\Psi}^{*}$ are not independent, making it difficult to analyze the probability model on $\boldsymbol{\Phi}$ directly. On the other hand, if the rows of $\boldsymbol{\Psi}$ are chosen uniformly at random with replacement from $\boldsymbol{\Psi}_{0}$, we can apply a matrix Bernstein inequality to estimate the operator norm of $\boldsymbol{\Psi}_{S}^{*}\boldsymbol{\Psi}_{S}-\textbf{I}_{S}$. We first state the following Lemma which follows almost exactly from Lemma 4.3 in [ling_strohmer_2015]. The only minor change in the proof is in the estimation of the bounds of $R$ and $\delta^{2}$ for the matrix Bernstein inequality.
@@
For any fixed matrix $\mathbb{B}\in\mathbb{C}^{N\times k}$ and $\boldsymbol{\Psi}^{*}\in\mathbb{C}^{N\times M}$ with its rows chosen uniformly at random with replacement from the rows of $\boldsymbol{\Psi}_{0}\in\mathbb{C}^{N\times M},\boldsymbol{\Psi}_{0}^{*}\boldsymbol{\Psi}_{0}=\mathbf{I}_{N}$. For the matrix $\boldsymbol{\Phi}$ defined in (24),
$\|{{\mathbf{\mathbf{\Phi}}}_{S}^{*}\mathbf{\mathbf{\Phi}}_{S}-\mathbf{I}_{S}}\|_{F\rightarrow F}\leq\delta$ (40)
@@
The inequality holds because whenever k is smaller than m as, $T_{k}\subseteq T_{m}$ Hence, $\mathbf{\mathbf{\Phi}}_{T_{k}}^{*}\mathbf{\mathbf{\Phi}}_{T_{k}}\preccurlyeq\mathbf{\mathbf{\Phi}}_{T_{m}}^{*}\mathbf{\mathbf{\Phi}}_{T_{m}}$.

### v3

@@ 4.2 A sufficient condition
The derivation of the lemma is a direct application of Lemma 7 of [4] which relies on the sub-differential of $\partial(\|\cdot\|_{(1,2)})$. Given $\textbf{Z}\in\mathbb{C}^{kN\times C}$, the sub-differential $\partial_{\textbf{Z}}(\|\cdot\|_{(1,2)})$ is the set
$\{\textbf{V}\in\mathbb{C}^{kN\times C}|\textbf{V}_{T_{i}}:=\frac{\textbf{Z}_{T_{i}\times[C]}}{\|\textbf{Z}_{T_{i}\times[C]}\|_{F}},\text{if }\|\textbf{Z}_{T_{i}\times[C]}\|_{F}\neq 0,\|\textbf{V}_{T_{i}}\|_{F}\leq 1,\text{otherwise}\}.$
In the following sections, we will prove that with the assumptions we have made, the parameters $\delta$ and $\beta$ will be small with high probability. We will also construct a dual certificate $\boldsymbol{\mathcal{Y}}$ with small $\eta$, $\theta$ and $\tau$. Most of the proof procedure follows well-established methods in the compressive sensing literature, such as those described in [1] and [4].
@@ 4.3 Conditioning of $\mathbf{A}_{S}$
Next, we give the conditions on when our measurement matrix $\mathbf{A}$ satisfies the condition $\|\textbf{A}_{S}^{*}\textbf{A}_{S}-\textbf{I}_{S}\|_{2\rightarrow 2}\leq\delta$. Recall our forward matrix $\mathbf{A}_{S}=\mathbf{\mathbf{F}}_{\Omega}\mathbf{\Phi}_{S}$, the challenges come from the fact that its rows are dependent. Instead of bounding the norm directly, we apply the triangular inequality and work on bounding the two terms in $\underbrace{\|\mathbf{\Phi}_{S}^{*}\mathbf{\textbf{F}}_{\Omega}^{*}\mathbf{\textbf{F}}_{\Omega}\mathbf{\Phi}_{S}-\mathbf{\Phi}_{S}^{*}\mathbf{\Phi}_{S}\|_{2\rightarrow 2}}_{(1)}+\underbrace{\|\mathbf{\Phi}_{S}^{*}\mathbf{\Phi}_{S}-\mathbf{I}_{S}\|_{2\rightarrow 2}}_{(2)}$ separately. Indeed, even when the full Fourier space is sampled, i.e., $\mathbf{F}_{\Omega}=\mathbf{F}$, the first term (1) is zero and we would at the very least, require the condition (2) to be satisfied.
@@
$\|\mathbf{\Phi}_{S}^{*}\mathbf{\Phi}_{S}-\textbf{I}_{S}\|_{2\rightarrow 2}\leq\eta$ (35)
@@
The probability is over the randomness in B. The proof of the Lemma is derived from Lemma 4.3 of [1], where the authors consider the case B to be fixed but the rows of $\mathbf{\Psi}_{S}$ are chosen uniformly at random with replacement from the DFT matrix.
@@
The next lemma is concerned with the tail bound for $\|\mathbf{\Phi}_{S}^{*}\mathbf{\textbf{F}}_{\Omega}^{*}\mathbf{\textbf{F}}_{\Omega}\mathbf{\Phi}_{S}-\mathbf{\Phi}_{S}^{*}\mathbf{\Phi}_{S}\|_{2\rightarrow 2}$.
@@
Suppose that the rows of $\mathbf{F}_{\Omega}$ are chosen uniformly at random without replacement from $\mathbf{F}$, where $\mathbf{F^{*}F}=\mathbf{F}\mathbf{F}^{*}=\frac{N}{L}$. For a realization of $\mathbf{\Phi}_{S}$ with $\|\mathbf{\Phi}_{S}^{*}\mathbf{\Phi}_{S}-\mathbf{I}_{S}\|_{2\rightarrow 2}\leq\eta_{1}$,
$\|\mathbf{\Phi}_{S}^{*}\textbf{F}_{\Omega}^{*}\textbf{F}_{\Omega}\mathbf{\Phi}_{S}-\mathbf{\Phi}_{S}^{*}\mathbf{\Phi}_{S}\|_{2\rightarrow 2}\leq\eta,\,\text{with probability at least}\,\,1-\epsilon,$
provided that
@@ 4.4 Dual Certificate
Suppose that either each column of H is chosen independently and uniformly at random with replacement from the columns of an orthonormal basis $\textbf{W}\in\mathbb{C}^{k\times k}$ or that each column of H is chosen independently and uniformly at random from the complex sphere $S_{\mathbb{C}}^{k-1}$ (i.e., $h_{l}\in\mathbb{C}^{k}\text{ with }\|h_{l}\|=1$). Conditioned on $\|\textbf{A}_{S}^{*}\textbf{A}_{S}-\textbf{I}\|_{2\rightarrow 2}\leq\eta$, if
@@
Equivalently, we can also state the result in Lemma 8 as $\max\limits_{j\in[|S^{c}|]}\|\mathcal{P}_{S^{c}_{j}}\boldsymbol{\mathcal{Y}}\|_{F}<1/2$ with probability at least $1-N^{-\alpha}$ if
@@
where $\tilde{C}_{\alpha}$ grows linearly with respect to $\alpha$. This is the same form we adopt in the statement of the main theorem.
Remark: Note that increasing the number of channels $C$ increases the probability of recovery. And when $C$ is sufficiently large, the failure probability for recovery is dominated by the other sufficient conditions required in Lemma 3 hence further increasing $C$ does not provide any additional benefits in terms of the probability of recovery. Furthermore, the lower bound on $L$ in Lemma 8 is automatically met due to the bound $L\sim O(kn\log(k^{2}n^{2}N))$ induced by the condition $\|\textbf{A}_{S}^{*}\textbf{A}_{S}-\textbf{I}\|_{2\rightarrow 2}\leq\eta$ from Lemma 5.
@@ 5.1 Empirical success rate for the proposed method
For the first set of experiments, we follow the same setup as in Ling and Strohmer [1] and calculate the empirical probability of success of the proposed method. Fix the number of measurements $L=128$ and the underlying signal dimension $N=256$, and solve the optimization problem (28) for varying $(k,n),k,n\in[15]$ and $C\in\{1,2,4\}$. For each $C$ and $(k,n)$, we generate $\textbf{z}\in\mathbb{C}^{N}$ with random support with cardinality $n$ and the nonzero entries of z and $\textbf{H}\in\mathbb{C}^{k\times C}$ drawn from $\mathcal{N}(0,1)$. Each time, we draw $L$ points uniformly from the Fourier space. We repeat the experiment ten times, and each time new H and z are generated.
We count an experiment as a success if the relative error between the solution of our optimization problem $\hat{\textbf{X}}$ and the truth $\textbf{X}_{0}$ is less than $0.01\%$, i.e., $\frac{\|\hat{\textbf{X}}-\textbf{X}_{0}\|_{F}}{\|\textbf{X}_{0}\|_{F}}\leq 10^{-4}$. And the empirical success rate is a fraction defined as (total number of successes) / (total number of experiments). We use the cvxpy [63, 64] package of Python with the MOSEK [65] solver to solve the mixed norm optimization problem (28).
@@
Next, we illustrate the barrier for reconstruction with the proposed framework when $\boldsymbol{\Psi}$ is a Wavelet transform. From the theory, a small mutual coherence, i.e., $\mu^{\Psi}_{max}\sim O(1)$, is required for the performance of the compressive sensing method. However, the Daubechies wavelet, which can be used to sparsify MRI images [33], has high coherence. We followed the non-uniform sampling scheme in the Fourier measurements as suggested by [33, 66] as an attempt to improve the results. However, even when $L=N$ (when Fourier space is fully sampled), recovery is poor. We use the Daubechies 4 wavelet as an illustration, and the results are shown in Figure 4.
@@ 5.1.1 Improved results compared to $\ell_{1}$ minimization
5.1.2 Minimal L required for exact recovery is proportional to knkn
@@ 5.2 Reconstruction of the analytical phantom
We use the analytical phantom [7] of size $256$ by $256$ and randomly generated coil sensitivities following the Biot–Savart law. It is important to note that we do not manually enforce sparsity in our signal. For example, one may manually use a sparse estimation of the phantom x, e.g., by crafting the signal by thresholding $\textbf{z}=\boldsymbol{\Psi}\textbf{x}=0,\text{ if }|\boldsymbol{\Psi}\textbf{x}|<\tau$ and set $\textbf{x}=\boldsymbol{\Psi}^{*}\textbf{z}$. The thresholding effectively reduces the dimensionality of the problem and would favor the reconstruction results. However, such manipulation is only possible in simulations when the true signal is known beforehand. Our approach is more flexible and allows for sparsity deficiency in the underlying signal.
@@
To assess the performance of the proposed method with respect to the number of coils, we plot the average reconstruction error versus the number of measurements per coil, for different numbers of coils and levels of Gaussian noise. We adopted a similar setup as that in [50, 51] where the multi-coil signal was generated by multiplying a small analytical phantom (N = 529) by the coil sensitivities (k=6) and optionally adding Gaussian noise. The average reconstruction error (relative $\ell_{2}$ distance) over 10 different random Fourier space undersampling realizations was computed for different numbers of measurements and coil elements. A DCT transform was employed to sparsify the image and no strict sparsity was enforced in the signal. The number of required measurements to achieve the noise floor reconstruction decreases with the number of coils for both the noise-free and the noisy case. These trends are consistent with previous numerical experiments for the known calibration pMRI reconstruction, e.g. [50]. The plots confirm that with a large number of coils, the number of measurements per coil sufficient to achieve a baseline level of reconstruction error is significantly smaller than the true dimensionality of the signal.
@@ 5.3 Comparison with the ENLIVE method
Next, we compare our reconstruction results with the ENLIVE formulation (17) in the paper by Holme et al. [47]. We reproduced the reconstructed brain and knee images exactly as shown in the original paper using the data and source code provided in https://github.com/mrirecon/enlive. We then compared these results with our own reconstructions. Figures 11 and 12 show calibrationless variable-density Poisson-disc undersampled reconstructions with differing undersampling factors comparing ENLIVE to our proposed method. Figure 13 shows the reconstructions of the simulated analytical phantom from uniformly subsampled Fourier space with different undersampling factors. As a reference, the corresponding sampling patterns are shown in the third columns of each of the plots.
@@ 6.1 Proof of Lemma 4
Recall that $\boldsymbol{\Phi}=\begin{bmatrix}(\bf{\bar{q}}_{1}\bigotimes\bf{b}_{1})^{*}\\ \vdots\\ (\bf{\bar{q}}_{N}\bigotimes\bf{b}_{N})^{*}\end{bmatrix}$, where each row of $\boldsymbol{\Phi}$ depends only on one row of B and $\boldsymbol{\Psi}^{*}$. Due to symmetry, the probability distribution of $\mathbb{\Phi}$ is unchanged if we assume B is fixed but the rows of $\mathbf{\Psi}^{*}$ are random. More precisely, we assume the rows of $\mathbf{\Psi}^{*}\in\mathbb{C}^{N\times N}$ are chosen uniformly at random without replacement from some $\mathbf{\Psi}_{0}^{*}\in\mathbb{C}^{N\times N}$ with $\mathbf{\Psi}_{0}^{*}\mathbf{\Psi}_{0}=\textbf{I}_{N}$. Under such a model, the rows of $\mathbf{\Psi}^{*}$ are not independent, making it difficult to analyze the probability model on $\boldsymbol{\Phi}$ directly. On the other hand, if the rows of $\boldsymbol{\Psi}$ are chosen uniformly at random with replacement from $\boldsymbol{\Psi}_{0}$, we can apply a matrix Bernstein inequality to estimate the operator norm of $\boldsymbol{\Psi}_{S}^{*}\boldsymbol{\Psi}_{S}-\textbf{I}_{S}$. We first state the following Lemma which follows almost exactly from Lemma 4.3 in [1]. The only minor change in the proof is in the estimation of the bounds of $R$ and $\delta^{2}$ for the matrix Bernstein inequality.
@@
 [^3] In our application, we use $M=N$.
For any fixed matrix $\mathbb{B}\in\mathbb{C}^{N\times k}$ and $\boldsymbol{\Psi}^{*}\in\mathbb{C}^{N\times M}$ with its rows chosen uniformly at random with replacement from the rows of $\boldsymbol{\Psi}^{*}_{0}\in\mathbb{C}^{N\times M},\boldsymbol{\Psi}_{0}^{*}\boldsymbol{\Psi}_{0}=\mathbf{I}_{N}$. For the matrix $\boldsymbol{\Phi}$ defined in (24),
$\|{{\mathbf{\mathbf{\Phi}}}_{S}^{*}\mathbf{\mathbf{\Phi}}_{S}-\mathbf{I}_{S}}\|_{2\rightarrow 2}\leq\delta$ (40)
@@
The inequality holds because whenever $T\subseteq\hat{T}$, we have $\mathbf{\mathbf{\Phi}}_{T}^{*}\mathbf{\mathbf{\Phi}}_{T}\preccurlyeq\mathbf{\mathbf{\Phi}}_{\hat{T}}^{*}\mathbf{\mathbf{\Phi}}_{\hat{T}}$.